# Quaternion-Based Sliding Mode Control for Six Degrees of Freedom Flight Control of Quadrotors

### v2

@@ I Introduction
Quadrotors require a robust controller to maintain stability and precise and agile maneuverability in the face of model uncertainties and external disturbances. One common control methodology that comes to mind when discussing robust control is sliding mode control (SMC), known for its simplicity and robustness [1]. SMC is proven effective in many application areas [2, 3, 4, 5]; however, its current implementations for quadrotors have certain shortcomings.
Previous work on sliding mode controller design for quadrotors can be categorized based on their formulation in modeling the vehicle rotational dynamics. An ample body of literature relies on the Euler angle representation of the vehicle attitude [6, 7, 8, 9, 10]. While robust 6-DOF flight control can be achieved with such methods, they are only suitable for small roll- or pitch-angle flight.
@@
The above equations have been the basis of numerous SMC designs for quadrotors [6, 7, 8, 9, 10]. While their particular second-order structure lends well to the SMC formulations, they are only valid for small $\phi$ and $\theta$ values. Thus, as we will show in our results, the controllers using such models struggle with maneuvers that entail large pitch and roll angles.
Another category of related papers relies on a quaternion-based representation of attitude. Several papers in this category lack 6-DOF SMC designs, applying SMC for only position control or attitude control [11, 12]. Several other papers have explored 6-DOF SMC designs [13, 14, 15]; however, they suffer from one or two of the following limitations: (i) the control architecture is complex, and (ii) there is the quaternion unwinding problem that can lead to longer-than-necessary attitude maneuvers.
The third category of papers uses the geometric control approaches, designing SMC in $\mathrm{SE}(3)$ [16, 17, 18]. In this approach, the attitude-tracking error is defined as ${\mathbf{R}}_{e}={\mathbf{R}}^{\top}_{d}\mathbf{R}$, where $\mathbf{R}$ is the rotation matrix, and the indices $e$ and $d$ represent the error and desired value [19]. This method is prone to slow convergence [20, 21] and leads to almost global stability [22].
In this paper, we address the above challenges by developing a new quaternion-based 6-DOF sliding mode controller, avoiding the aforementioned limitations of Euler-based and $\mathrm{SE}(3)$ approaches. We adopt the method presented in [21] to eliminate the unwinding problem of quaternions. We apply a cascaded control design, where the position controller output in the outer loop is used to generate the desired trajectory for the attitude controller. We use a coordinate-free desired attitude generation method [19] with no simplification of rotational dynamics, enabling effective flight control during aggressive maneuvers in the presence of model uncertainties and external disturbances. Our controller features global stability and uses the inherent characteristics of quaternion dynamics in $\mathbb{S}^{3}$ to achieve such remarkable performance. We compare our results with common controllers [19, 23, 22], and show consistent improvements in flight performance while minimizing actuator effort and saturation.
@@ II Quadrotor dynamics
$\mathbf{G}=\begin{bmatrix}1&1&1&1\\ l\;\sin(\beta)&-l\;\sin(\beta)&-l\;\sin(\beta)&l\;\sin(\beta)\\ -l\;\cos(\beta)&l\;\cos(\beta)&-l\;\cos(\beta)&l\;\sin(\beta)\\ -c_{q}/c_{t}&-c_{q}/c_{t}&c_{q}/c_{t}&c_{q}/c_{t}\end{bmatrix}$, (7)
@@ III-B Trajectory generation for the attitude controller
The auxiliary variable $\boldsymbol{\kappa}$ can be used to generate the desired quaternion ${\mathbf{q}}_{d}$, desired angular velocity $\boldsymbol{\omega}_{d}$, and desired angular acceleration $\boldsymbol{\alpha}_{d}$ for the attitude controller. Inspired by the method in [19], we consider the desired rotation matrix $\mathbf{R}_{d}=\left[\mathbf{b}_{1_{d}},\mathbf{b}_{2_{d}},\mathbf{b}_{3_{d}}\right]$, where $\mathbf{b}_{3_{d}}={\boldsymbol{\kappa}}/{\lVert\boldsymbol{\kappa}\rVert}$. The goal is to find $\mathbf{b}_{1_{d}}$ and $\mathbf{b}_{2_{d}}$ to construct ${\mathbf{R}}_{d}$, which will lead to finding ${\mathbf{q}}_{d}$, $\boldsymbol{\omega}_{d}$, and $\boldsymbol{\alpha}_{d}$.
@@
To generate $\boldsymbol{\omega}_{d}$ and $\boldsymbol{\alpha}_{d}$, we use $\dot{\mathbf{R}}_{d}=[\dot{\mathbf{b}}_{1_{d}},\dot{\mathbf{b}}_{2_{d}},\dot{\mathbf{b}}_{3_{d}}]$ and $\ddot{\mathbf{R}}_{d}=[\ddot{\mathbf{b}}_{1_{d}},\ddot{\mathbf{b}}_{2_{d}},\ddot{\mathbf{b}}_{3_{d}}]$. According to [19], we have
@@ III-C Attitude controller
For the attitude controller, we adopt the method proposed in [21]. The advantage of this approach is capturing the geometry of $\mathbb{S}^{3}$ and global stability as opposed to common geometric methods that are almost globally stable [22]. To this end, let us define the quaternion and angular velocity errors as
@@
and $\boldsymbol{\lambda}_{q}$ is a design parameter with positive elements. The term $\operatorname{sgn_{+}}(q_{w_{e}})$ is the key in this design. It maintains the control input continuity, thus eliminating the unwinding behavior of quaternion-based controllers. Next, according to [21], the attitude control law becomes
@@ III-D Stability analysis
provided that the design parameters $\mathbf{K}_{\xi}$ and $\mathbf{K}_{q}$ are sufficiently large. The steps to show the stability are straightforward using standard SMC stability analysis [1], and the mathematical treatment presented in [21]. To avoid repetitions, and use the limited space to present our results, we skip the details of this part.
@@ IV RESULTS
This section presents our simulation results. We compare our method with two common 6-DOF flight control techniques, the geometric controller [19], and the Euler-based sliding mode controller [24]. Also, to highlight the advantages of our attitude controller, we compare our results with a commonly used quaternion-based proportional-derivative (PD) attitude controller [23] combined with our position controller. The simulation scenarios consist of (1) flip-maneuver, and (2) lemniscate trajectory tracking, both in the presence of parametric uncertainties and external disturbances. Table I gives the parameter values used in simulations. Note that we meticulously fine-tuned all controllers over many trials to ensure their best performance.
@@
d) Geometric controller parameters [19]:
@@
e) Euler-based SMC parameters [24]:
@@
f) Quaternion-based PD controller parameters [23]:

### v3

@@ I Introduction
Quadrotors require a robust controller to maintain stability and precise and agile maneuverability in the face of model uncertainties and external disturbances. One common control methodology that comes to mind when discussing robust control is sliding mode control (SMC), known for its simplicity and robustness [slotine1991applied]. SMC is proven effective in many application areas [DANGELO2024104554, nemati2019fast, 10156565, faieghi2012novel]; however, its current implementations for quadrotors have certain shortcomings.
Previous work on sliding mode controller design for quadrotors can be categorized based on their formulation in modeling the vehicle rotational dynamics. An ample body of literature relies on the Euler angle representation of the vehicle attitude [xu2006sliding, zheng2014second, xiong2017global, izadi2024high, thanh2018quadcopter]. While robust 6-DOF flight control can be achieved with such methods, they are only suitable for small roll- or pitch-angle flight.
@@
The above equations have been the basis of numerous SMC designs for quadrotors [xu2006sliding, zheng2014second, xiong2017global, izadi2024high, thanh2018quadcopter]. While their particular second-order structure lends well to the SMC formulations, they are only valid for small $\phi$ and $\theta$ values. Thus, as we will show in our results, the controllers using such models struggle with maneuvers that entail large pitch and roll angles.
Another category of related papers relies on a quaternion-based representation of attitude. Several papers in this category lack 6-DOF SMC designs, applying SMC for only position control or attitude control [esmail2022attitude, abaunza2019quadrotor]. Several other papers have explored 6-DOF SMC designs [serrano2023terminal, arellano2015quaternion, sanwale2020quaternion]; however, they suffer from one or two of the following limitations: (i) the control architecture is complex, and (ii) there is the quaternion unwinding problem that can lead to longer-than-necessary attitude maneuvers.
The third category of papers uses the geometric control approaches, designing SMC in $\mathrm{SE}(3)$ [garcia2020robust, gong2020adaptive, ren2023adaptive]. In this approach, the attitude-tracking error is defined as ${\mathbf{R}}_{e}={\mathbf{R}}^{\top}_{d}\mathbf{R}$, where $\mathbf{R}$ is the rotation matrix, and the indices $e$ and $d$ represent the error and desired value [lee2010geometric]. This method is prone to slow convergence [teng2022lie, lopez2020sliding] and leads to almost global stability [bhat2000topological].
In this paper, we address the above challenges by developing a new quaternion-based 6-DOF sliding mode controller, avoiding the aforementioned limitations of Euler-based and $\mathrm{SE}(3)$ approaches. We adopt the method presented in [lopez2020sliding] to eliminate the unwinding problem of quaternions. We apply a cascaded control design, where the position controller output in the outer loop is used to generate the desired trajectory for the attitude controller. We use a coordinate-free desired attitude generation method [lee2010geometric] with no simplification of rotational dynamics, enabling effective flight control during aggressive maneuvers in the presence of model uncertainties and external disturbances. Our controller features global stability and uses the inherent characteristics of quaternion dynamics in $\mathbb{S}^{3}$ to achieve such remarkable performance. We compare our results with common controllers [lee2010geometric, fresk2013full, bhat2000topological], and show consistent improvements in flight performance while minimizing actuator effort and saturation.
@@ II Quadrotor dynamics
$\mathbf{G}=\begin{bmatrix}1&1&1&1\\ -l\;\sin(\beta)&l\;\sin(\beta)&l\;\sin(\beta)&-l\;\sin(\beta)\\ l\;\cos(\beta)&-l\;\cos(\beta)&l\;\cos(\beta)&-l\;\cos(\beta)\\ -c_{q}/c_{t}&-c_{q}/c_{t}&c_{q}/c_{t}&c_{q}/c_{t}\end{bmatrix}$. (7)
@@ III-B Trajectory generation for the attitude controller
The auxiliary variable $\boldsymbol{\kappa}$ can be used to generate the desired quaternion ${\mathbf{q}}_{d}$, desired angular velocity $\boldsymbol{\omega}_{d}$, and desired angular acceleration $\boldsymbol{\alpha}_{d}$ for the attitude controller. Inspired by the method in [lee2010geometric], we consider the desired rotation matrix $\mathbf{R}_{d}=\left[\mathbf{b}_{1_{d}},\mathbf{b}_{2_{d}},\mathbf{b}_{3_{d}}\right]$, where $\mathbf{b}_{3_{d}}={\boldsymbol{\kappa}}/{\lVert\boldsymbol{\kappa}\rVert}$. The goal is to find $\mathbf{b}_{1_{d}}$ and $\mathbf{b}_{2_{d}}$ to construct ${\mathbf{R}}_{d}$, which will lead to finding ${\mathbf{q}}_{d}$, $\boldsymbol{\omega}_{d}$, and $\boldsymbol{\alpha}_{d}$.
@@
To generate $\boldsymbol{\omega}_{d}$ and $\boldsymbol{\alpha}_{d}$, we use $\dot{\mathbf{R}}_{d}=[\dot{\mathbf{b}}_{1_{d}},\dot{\mathbf{b}}_{2_{d}},\dot{\mathbf{b}}_{3_{d}}]$ and $\ddot{\mathbf{R}}_{d}=[\ddot{\mathbf{b}}_{1_{d}},\ddot{\mathbf{b}}_{2_{d}},\ddot{\mathbf{b}}_{3_{d}}]$. According to [lee2010geometric], we have
@@ III-C Attitude controller
For the attitude controller, we adopt the method proposed in [lopez2020sliding]. The advantage of this approach is capturing the geometry of $\mathbb{S}^{3}$ and global stability as opposed to common geometric methods that are almost globally stable [bhat2000topological]. To this end, let us define the quaternion and angular velocity errors as
@@
and $\boldsymbol{\lambda}_{q}$ is a design parameter with positive elements. The term $\operatorname{sgn_{+}}(q_{w_{e}})$ is the key in this design. It maintains the control input continuity, thus eliminating the unwinding behavior of quaternion-based controllers. Next, according to [lopez2020sliding], the attitude control law becomes
@@ III-D Stability analysis
provided that the design parameters $\mathbf{K}_{\xi}$ and $\mathbf{K}_{q}$ are sufficiently large. The steps to show the stability are straightforward using standard SMC stability analysis [slotine1991applied], and the mathematical treatment presented in [lopez2020sliding]. To avoid repetitions, and use the limited space to present our results, we skip the details of this part.
@@ IV RESULTS
This section presents our simulation results. We compare our method with two common 6-DOF flight control techniques, the geometric controller [lee2010geometric], and the Euler-based sliding mode controller [nadda_adaptive_2018]. Also, to highlight the advantages of our attitude controller, we compare our results with a commonly used quaternion-based proportional-derivative (PD) attitude controller [fresk2013full] combined with our position controller. The simulation scenarios consist of (1) flip-maneuver, and (2) lemniscate trajectory tracking, both in the presence of parametric uncertainties and external disturbances. Table I gives the parameter values used in simulations. Note that we meticulously fine-tuned all controllers over many trials to ensure their best performance.
@@
d) Geometric controller parameters [lee2010geometric]:
@@
e) Euler-based SMC parameters [nadda_adaptive_2018]:
@@
f) Quaternion-based PD controller parameters [fresk2013full]: